A copper wire and a steel wire having the same cross-section area are fastened end to end stretched by a force F. the lengths of copper and steel wires are in the ratio of 2:1 and their moduli of elasticity are in the ratio of 1:2. What is the ratio of their extensions?
(a) 1:2 (b) 4:1
(c) 2:1 (d) 1:4

Grade:12

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

The answer is option b

Regards
Vikas TU
3 years ago
e=MgL/πr^2Y
∴ec/es=Lc/Ls×Ys/Yc=2/1×2/1=4/1=4:1
Good Luck …............................................................